“A horizontal spring-mass system is consists of a mass sliding on a frictionless surface with negligible friction. When the block has velocity v = +vmax/2, what fraction of the total energy is the elastic potential energy?”
Answer #1

TE = P.E+KE + Kw = mare - 0 I kn = dk (24) + 2 ml khoe) use eq 0 Im Wear) - I m non)) = 1 k.) & meron) { } } = 1 ko : f eac

Please rate the work if you are satisfied with the solution or else kindly ask your doubts or queries in the comment section below